# Recent advances in text embedding: A Comprehensive Review of Top-Performing Methods on the MTEB Benchmark

### v1

@@ 3 Data focused universal text embeddings
One way to learn the universal text embedding is using a multi-stage contrastive learning strategy with improved training data mixture in terms of data quantity, quality and diversity as summarized in Table 1. For example, GTE (?) uses diverse datasets for both pre-training and fine-tuning stage. BGE (?) introduces a compressive data package C-Pack, while E5 (?) constructed a curated web-scale text pair dataset named Colossal Clean text Pairs (CCPairs) containing heterogeneous training signals by combining various semi-structured data sources along with aggressive filtering (270M text pairs filtered from 1.3B noisy text pairs) with a consistency-based filter to improve data quality (?). Some works like GISTEmbed (?) also focus on improving the quality of hard negatives used for training. More details about each text embedding methods can be found below.
@@
GTE (?) Substantial performance gains are achieved by notably augmenting data volume during both unsupervised pre-training (800M text pairs used for pre-training) and supervised fine-tuning stages with diverse mixture of datasets from multiple sources.
BGE (?) The largest dataset C-MTP was developed for general Chinese embedding with the focuses on: 1. data quality improvement by filtering the irrelevant text pairs in unlabelled data for general purpose fine-tuning (around 100M text pairs); 2. multi-task high quality labelled data (838,465 text pairs) for task-specific fine-tuning. Note: English data (for English version of BGE) is 2 times larger than the Chinese data.
GISTEmbed (?) GIST is fine-tuned on top of BGE on MEDI and MTEB classification datasets with improved in-batch negative data quality.
E5 (?) Development of CCPairs: curated large-scale text pair dataset by harvesting heterogeneous semi-structured data sources using consistency-based filter for quality improvement (reducing 1.3B text pairs to 270M text pairs for pre-training).
Multilingual-E5 (?) Multilingual focus: diverse mixture of multilingual text pairs obtained from various sources (1B text pairs). Additional 500k synthetic data generated by GPT-3.5/4 which encompasses 150k unique instructions and covers 93 languages were used for fine-tuning.
@@
With the focus on developing a unified more comprehensive model for general text representation to address a multitude of downstream tasks, the authors from (?) introduce a multi-stage contrastive learning strategy with diverse training data mixture: in the initial stage, a large corpus of open-source data without any filtering or cleaning are used to learn basic language patterns with unsupervised contrastive learning; in the second stage, supervised fine-tuning refines the embeddings using contrastive learning with a smaller, high-quality dataset. At both stages, the number of training data are significantly increased.
For a query $q$, a relevant/positive document $d^{+}$, a set of irrelevant/negative documents $D_{-}=\{d_{-}^{1},...,d_{-}^{n}\}$, the InfoNCE loss (?) is defined as in Equation 1:
@@
In GTE, given a batch of positive text pair samples $\{(q_{1},d_{1}),(q_{2},d_{2}),...,(q_{n},d_{n})\}$, the authors propose an improved contrastive loss (icl) can be viewed as a combination of loss variants proposed by (?, ?, ?):
@@
    • Web Search: MS MARCO (?) passage retrieval benchmarks where hard negatives are mined by sampling from high-ranked documents retrieval system, excluding positive ones.
    • Open QA: Natural Questions (NQ), Trivia QA (?, ?), Web Questions, HotpotQA (?), etc. Top ranked passage by retrieval system which do not include answer to the question is regarded as hard negatives.
    • Natural Language Inference: MNLI (?) and SNLI (?), entailment as positive pairs and contradiction as negative pairs
    • Fact Verification: training set from FEVER (?)
    • Paraphrase: Quora (?) and StackExchange Dupquestion
    • Others: miscellaneous datasets from different NLP tasks and domains released in MEDI (?) and BERRI (?).
@@
Similar to the objective of GTE, BGE also tries to learn general-purpose text embeddings, a comprehensive, unified embedding model which is capable of managing all types of uses, including retrieval, ranking, and classification, across various application settings such as question answering, language modeling, and conversation (?). BGE introduces C-Pack, a comprehensive package designed to advance the general Chinese embedding (other languages version of BGE are also available), along with their training recipe: pre-training of an embedding-oriented text encoder, general-purpose contrastive learning, and task-specific fine-tuning. BERT-like architecture is used by BGE models where the last layer’s hidden state of the special token [CLS] is trained to work as the embedding (unlike GTE). Another major difference from GTE is that BGE uses instruction-based fine-tuning to deal with potentially mutually contradicted tasks: a task specific instruction which describes the nature of the task (e.g. search relevant passages for the query) is added to the query side for each text pair. Some other details about BGE include:
  • Pre-training data (English version): unsupervised datasets including datasets like Wikipedia, CC-net, StackExchange, Reddit, S2ORC (?), and datasets from sentence-transformers [^2] https://huggingface.co/datasets/sentence-transformers/embedding-training-data.
  • Fine-tuning data (English version): supervised datasets including NLI (?), FEVER (?), NQ (?, ?), HotpotQA (?), Quora (?), StackExchange Duplicates and MEDI (?).
@@
    • Pre-training: purely rely on in-batch negative samples (?) and resort to a big batch size (as large as 19,200) to improve the discriminativeness of the embedding.
    • Fine-tuning: in addition to the in-batch negative samples, one hard negative sample is mined for each text pair from the task’s original corpus, following the ANN-style sampling strategy in (?)
@@
Guided In-sample Selection of Training Negatives for Text Embedding Fine-tuning (GISTEmbed) GIST-large-Embedding-v0 is another top performing text embeddings on the MTEB benchmark which uses $BGE_{large}$ as backbone. The main focus of GISTEmbed is to propose a novel strategy that enhances in-batch negative selection during contrastive training through a guide model (?), which improves the baseline performance slightly. However, the GIST-large-Embedding-v0 performance increase on MTEB benchmark compared to $BGE_{large}$ is limited (0.11%). It is difficult to analyze if the limited performance increase is due to the proposed guided in-sample negative selection or due to the fact that they added in-domain MTEB training data to fine-tune the BGE embedding models.
@@
With the objective of creating high-quality general-purpose text embeddings suitable for any tasks requiring single-vector representations in both zero-shot or fine-tuned settings, the authors from (?) constructed a curated web-scale text pair dataset named Colossal Clean text Pairs (CCPairs) containing heterogeneous training signals by combining various semi-structured data sources such as CommunityQA, Common Crawl and Scientific papers along with aggressive filtering (270M text pairs filtered from 1.3B noisy text pairs) with a consistency-based filter to improve data quality (?). Some other details about E5 include:
  • Pre-training data: (post, comment) pairs from Reddit [^3] https://files.pushshift.io/reddit/, (question, upvoted answer) pairs from Stackexchange [^4] https://archive.org/details/stackexchange, (entity name + section title, passage) pairs from English Wikipedia, (title, abstract) and citation pairs from Scientific papers (?), and (title, passage) pairs from Common Crawl web pages [^5] https://commoncrawl.org/, various News sources and others including “SimpleWiki”, “GooAQ”, “WikiHow”, “Yahoo Answers” [^6] https://huggingface.co/datasets/ sentence-transformers/embedding-training-data.
  • Fine-tuning data: Natural Language Inference (NLI (?)), MS-MARCO passage ranking dataset (?), and Natural Questions (NQ) dataset (?, ?)
@@
    • $E5_{small}$: 33M, initialized from MiniLM (?)
@@
In order to extend the English E5 models, the authors from (?) released Multilingual-E5 series models by using diverse mixture of multilingual text pairs obtained from various sources with around 1 billion text pairs. The English E5 model recipe is used for the training procedure, which involves contrastive pre-training on 1 billion multilingual text pairs and fine-tuning on a blend of labeled datasets, with the Multilingual-E5-large-instruct adopting the data mixture from (?) that includes an additional 500k synthetic data created by GPT-3.5/4 and encompasses 150k unique instructions across 93 languages. Similar to BGE, instructions data are used to better inform embedding models about the task at hand for Multilingual-E5-large-instruct model. Some other details about Multilingual-E5 include:
@@
    • Multilingual-E5-small: 118M (initialized from multi-lingual MiniLM (?)),
    • Multilingual-E5-base: 278M (initialized from xlm-roberta-base (?)),
    • Multilingual-E5-large: 560M (initialized from xlm-roberta-large (?))
@@ 4 Loss focused universal text embeddings
Contrastive learning with InfoNCE loss (Equation 1) is used by most of the state of the art universal text embedding models. Several loss variants have been proposed by (?, ?, ?) and the authors of GTE propose an improved contrastive loss (Equation 2) which combines these loss variants. As many existing text embedding works employed the cosine function in their training objective to measure the pairwise semantic similarity, the authors from UAE (?) point out that there is the gradient vanishing issue due to the saturation zones of cosine function, which hinder the ability to learn subtle distinctions between texts in back propagation. Hence they propose a novel angle-optimized text embedding model called AnglE with angle optimization in a complex space which substantially improve the text embedding quality in various scenarios. Matryoshka Representation Learning (MRL) (?) and 2D Matryoshka Sentence Embeddings (2DMSE) propose new loss functions in order to reduce the computational cost of downstream tasks. More details about different novel losses proposed by the MTEB top performing universal text embedding models can be found below.
@@
Similar to GTE and BGE, UAE also uses the pre-trained BERT model (uncased BERT base model with 110M parameters) as the backbone model (note: UAE_large_V1 uses roberta-large (?) as the default backbone). As many existing text embedding works employed the cosine function in their training objective to measure the pairwise semantic similarity, the authors from (?) point out that there is the gradient vanishing issue due to the saturation zones of cosine function (as shown in Figure 3), which hinder the ability to learn subtle distinctions between texts in backpropagation. To deal with the problem of vanishing gradients, a novel angle-optimized text embedding model called AnglE is proposed by introducing angle difference optimization in a complex space which substantially improve the text embedding quality in various scenarios. Given input text embedding pair $(E(q),E(d))$, the chunking strategy (?) is used to get their representations in the complex space $(\mathbf{z},\mathbf{w})$, followed by the angle difference $\Delta\theta_{qd}$ between $\mathbf{z}$ and $\mathbf{w}$. Then the angle loss can be defined as:
@@
where $sim(E(i),E(j))$ is the similarity between the embedding of $i$ and the embedding of $j$. The authors also propose LLM-supervised learning (use LLMs as data annotators to label the pseudo-supervised data) to effectively deal with the domain-supervised data scarcity problem (?). Some other details about UAE include:
  • Training data: the NLI datasets MNLI (?) and SNLI (?), and/or LLM supervised data
@@
Matryoshka Representation Learning (MRL) (?)
Deploying deep representation or text embedding involves two steps: a constant forward pass to compute the representation, and its use for downstream applications (?, ?). The computation costs for the second step rise with the embedding dimensionality, data size, and label space, which can exceed the feature computation cost for large scale systems (?, ?). The rigid nature of these representations requires high-dimensional embedding vectors for different tasks, even though varying resource and accuracy constraints call for flexibility (?). Given that we can’t predict the computational and statistical demands for each subsequent task, fixed-capacity representations/embeddings may not always be suitable and could either exceed or fall short of the task’s requirements. Could we create an adaptable representation that can adjust to a variety of downstream tasks with fluctuating computational resources?
MRL introduces a novel method for learning representations of data through a nested structure to induce flexibility in the learned representation, similar to Russian Matryoshka dolls, which encodes information at different granularities and allows a single embedding to adapt to the computational constraints of downstream tasks (?). The representation/embedding $z$ is a $d$ dimensional vector, $M=[m_{1},m_{2},...d]$ are the chosen dimensions which define different representation sizes. MRL makes each of the first $m$ dimensions $z_{1:m}$ to be independently capable of being a general purpose representation of the data point $x$. Given a labelled dataset $D=\{(x_{1},y_{1}),(x_{2},y_{2}),...(x_{N},y_{N})\}$ where $N$ is the datasize and $y_{i}$ is the label of data $x_{i}$, MRL uses standard empirical risk minimization to optimize multi-class classification loss for each nested dimension $m\in M$ using a separate linear classifier, parameterized by $\mathbf{W}^{(m)}$:
@@
where $\mathcal{L}$ is the multi-class softmax cross-entropy loss function, $F(\cdot;\theta_{F})$ is the deep neural network to get the representation/embedding $z$, $c_{m}$ is the importance scales. The authors also show that MRL extends seamlessly to web-scale datasets across vision, language, and vision + language. The experimental results show that MRL can be effectively used for large-scale adaptive classification and retrieval, providing similar accuracy to fixed-feature baseline with a significantly smaller representation size, and offering a more cost-effective and faster adaptive shortlisting and re-ranking system (?).
@@
Despite MRL’s enhanced efficiency, it still necessitates going through all transformer layers before obtaining the transformer based text embedding, leading to significant compute and memory consumption. This raises questions about the impact of the fixed number of transformer layers on representation quality and the feasibility of using intermediate layers for sentence representation. With the aim to enhance the flexibility and scalability of the original MRL’s sentence embedding learning, two-dimensional Matryoshka Sentence Embedding (2DMSE) is proposed in (?). 2DMSE uses $BERT_{base}$ as backbone to encode text data $x$:
@@
The auxiliary information A is utilized for loss computation, typically indicating positive or negative samples or providing ranking details (?). During the same training step, a shallower Transformer layer $l$ is randomly chosen following a uniform distribution $l\sim U(1,L-1)$, and its complete embedding vector is directly utilized for representation learning:
@@
Based on MRL and 2DMSE, several well performing text embeddings including mxbai-embed-large-v1 (335M) and mxbai-embed-2d-large-v1 (335M) are released in (?). However, the training details of these models are not documented.

### v2

@@ 3 Data focused universal text embeddings
One way to learn the universal text embedding is using a multi-stage contrastive learning strategy with improved training data mixture in terms of data quantity, quality and diversity as summarized in Table 1. For example, GTE [12] uses diverse datasets for both pre-training and fine-tuning stage. BGE [35] introduces a compressive data package C-Pack, while E5 [21] constructed a curated web-scale text pair dataset named Colossal Clean text Pairs (CCPairs) containing heterogeneous training signals by combining various semi-structured data sources along with aggressive filtering (270M text pairs filtered from 1.3B noisy text pairs) with a consistency-based filter to improve data quality [64]. Some works like GISTEmbed [47] also focus on improving the quality of hard negatives used for training. More details about each text embedding methods can be found below.
@@
GTE [12] Substantial performance gains are achieved by notably augmenting data volume during both unsupervised pre-training (800M text pairs used for pre-training) and supervised fine-tuning stages with diverse mixture of datasets from multiple sources.
BGE [35] The largest dataset C-MTP was developed for general Chinese embedding with the focuses on: 1. data quality improvement by filtering the irrelevant text pairs in unlabelled data for general purpose fine-tuning (around 100M text pairs); 2. multi-task high quality labelled data (838,465 text pairs) for task-specific fine-tuning. Note: English data (for English version of BGE) is 2 times larger than the Chinese data.
GISTEmbed [47] GIST is fine-tuned on top of BGE on MEDI and MTEB classification datasets with improved in-batch negative data quality.
E5 [21] Development of CCPairs: curated large-scale text pair dataset by harvesting heterogeneous semi-structured data sources using consistency-based filter for quality improvement (reducing 1.3B text pairs to 270M text pairs for pre-training).
Multilingual-E5 [38] Multilingual focus: diverse mixture of multilingual text pairs obtained from various sources (1B text pairs). Additional 500k synthetic data generated by GPT-3.5/4 which encompasses 150k unique instructions and covers 93 languages were used for fine-tuning.
@@
With the focus on developing a unified more comprehensive model for general text representation to address a multitude of downstream tasks, the authors from [12] introduce a multi-stage contrastive learning strategy with diverse training data mixture: in the initial stage, a large corpus of open-source data without any filtering or cleaning are used to learn basic language patterns with unsupervised contrastive learning; in the second stage, supervised fine-tuning refines the embeddings using contrastive learning with a smaller, high-quality dataset. At both stages, the number of training data are significantly increased.
For a query $q$, a relevant/positive document $d^{+}$, a set of irrelevant/negative documents $D_{-}=\{d_{-}^{1},...,d_{-}^{n}\}$, the InfoNCE loss [65] is defined as in Equation 1:
@@
In GTE, given a batch of positive text pair samples $\{(q_{1},d_{1}),(q_{2},d_{2}),...,(q_{n},d_{n})\}$, the authors propose an improved contrastive loss (icl) can be viewed as a combination of loss variants proposed by [66, 67, 68]:
@@
    • Web Search: MS MARCO [69] passage retrieval benchmarks where hard negatives are mined by sampling from high-ranked documents retrieval system, excluding positive ones.
    • Open QA: Natural Questions (NQ), Trivia QA [70, 71], Web Questions, HotpotQA [72], etc. Top ranked passage by retrieval system which do not include answer to the question is regarded as hard negatives.
    • Natural Language Inference: MNLI [73] and SNLI [74], entailment as positive pairs and contradiction as negative pairs
    • Fact Verification: training set from FEVER [75]
    • Paraphrase: Quora [76] and StackExchange Dupquestion
    • Others: miscellaneous datasets from different NLP tasks and domains released in MEDI [77] and BERRI [36].
@@
Similar to the objective of GTE, BGE also tries to learn general-purpose text embeddings, a comprehensive, unified embedding model which is capable of managing all types of uses, including retrieval, ranking, and classification, across various application settings such as question answering, language modeling, and conversation [35]. BGE introduces C-Pack, a comprehensive package designed to advance the general Chinese embedding (other languages version of BGE are also available), along with their training recipe: pre-training of an embedding-oriented text encoder, general-purpose contrastive learning, and task-specific fine-tuning. BERT-like architecture is used by BGE models where the last layer’s hidden state of the special token [CLS] is trained to work as the embedding (unlike GTE). Another major difference from GTE is that BGE uses instruction-based fine-tuning to deal with potentially mutually contradicted tasks: a task specific instruction which describes the nature of the task (e.g. search relevant passages for the query) is added to the query side for each text pair. Some other details about BGE include:
  • Pre-training data (English version): unsupervised datasets including datasets like Wikipedia, CC-net, StackExchange, Reddit, S2ORC [78], and datasets from sentence-transformers [^2] https://huggingface.co/datasets/sentence-transformers/embedding-training-data.
  • Fine-tuning data (English version): supervised datasets including NLI [57], FEVER [75], NQ [70, 71], HotpotQA [72], Quora [76], StackExchange Duplicates and MEDI [77].
@@
    • Pre-training: purely rely on in-batch negative samples [79] and resort to a big batch size (as large as 19,200) to improve the discriminativeness of the embedding.
    • Fine-tuning: in addition to the in-batch negative samples, one hard negative sample is mined for each text pair from the task’s original corpus, following the ANN-style sampling strategy in [80]
@@
Guided In-sample Selection of Training Negatives for Text Embedding Fine-tuning (GISTEmbed) GIST-large-Embedding-v0 is another top performing text embeddings on the MTEB benchmark which uses $BGE_{large}$ as backbone. The main focus of GISTEmbed is to propose a novel strategy that enhances in-batch negative selection during contrastive training through a guide model [47], which improves the baseline performance slightly. However, the GIST-large-Embedding-v0 performance increase on MTEB benchmark compared to $BGE_{large}$ is limited (0.11%). It is difficult to analyze if the limited performance increase is due to the proposed guided in-sample negative selection or due to the fact that they added in-domain MTEB training data to fine-tune the BGE embedding models.
@@
With the objective of creating high-quality general-purpose text embeddings suitable for any tasks requiring single-vector representations in both zero-shot or fine-tuned settings, the authors from [21] constructed a curated web-scale text pair dataset named Colossal Clean text Pairs (CCPairs) containing heterogeneous training signals by combining various semi-structured data sources such as CommunityQA, Common Crawl and Scientific papers along with aggressive filtering (270M text pairs filtered from 1.3B noisy text pairs) with a consistency-based filter to improve data quality [64]. Some other details about E5 include:
  • Pre-training data: (post, comment) pairs from Reddit [^3] https://files.pushshift.io/reddit/, (question, upvoted answer) pairs from Stackexchange [^4] https://archive.org/details/stackexchange, (entity name + section title, passage) pairs from English Wikipedia, (title, abstract) and citation pairs from Scientific papers [78], and (title, passage) pairs from Common Crawl web pages [^5] https://commoncrawl.org/, various News sources and others including “SimpleWiki”, “GooAQ”, “WikiHow”, “Yahoo Answers”  [^6] https://huggingface.co/datasets/sentence-transformers/embedding-training-data.
  • Fine-tuning data: Natural Language Inference (NLI [74]), MS-MARCO passage ranking dataset [69], and Natural Questions (NQ) dataset [70, 71]
@@
    • $E5_{small}$: 33M, initialized from MiniLM [81]
@@
In order to extend the English E5 models, the authors from [38] released Multilingual-E5 series models by using diverse mixture of multilingual text pairs obtained from various sources with around 1 billion text pairs. The English E5 model recipe is used for the training procedure, which involves contrastive pre-training on 1 billion multilingual text pairs and fine-tuning on a blend of labeled datasets, with the Multilingual-E5-large-instruct adopting the data mixture from [16] that includes an additional 500k synthetic data created by GPT-3.5/4 and encompasses 150k unique instructions across 93 languages. Similar to BGE, instructions data are used to better inform embedding models about the task at hand for Multilingual-E5-large-instruct model. Some other details about Multilingual-E5 include:
@@
    • Multilingual-E5-small: 118M (initialized from multi-lingual MiniLM [81]),
    • Multilingual-E5-base: 278M (initialized from xlm-roberta-base [82]),
    • Multilingual-E5-large: 560M (initialized from xlm-roberta-large [82])
@@ 4 Loss focused universal text embeddings
Contrastive learning with InfoNCE loss (Equation 1) is used by most of the state of the art universal text embedding models. Several loss variants have been proposed by [66, 67, 68] and the authors of GTE propose an improved contrastive loss (Equation 2) which combines these loss variants. As many existing text embedding works employed the cosine function in their training objective to measure the pairwise semantic similarity, the authors from UAE [50] point out that there is the gradient vanishing issue due to the saturation zones of cosine function, which hinder the ability to learn subtle distinctions between texts in back propagation. Hence they propose a novel angle-optimized text embedding model called AnglE with angle optimization in a complex space which substantially improve the text embedding quality in various scenarios. Matryoshka Representation Learning (MRL) [62] and 2D Matryoshka Sentence Embeddings (2DMSE) propose new loss functions in order to reduce the computational cost of downstream tasks. More details about different novel losses proposed by the MTEB top performing universal text embedding models can be found below.
@@
Similar to GTE and BGE, UAE also uses the pre-trained BERT model (uncased BERT base model with 110M parameters) as the backbone model (note: UAE_large_V1 uses roberta-large [83] as the default backbone). As many existing text embedding works employed the cosine function in their training objective to measure the pairwise semantic similarity, the authors from [50] point out that there is the gradient vanishing issue due to the saturation zones of cosine function (as shown in Figure 3), which hinder the ability to learn subtle distinctions between texts in backpropagation. To deal with the problem of vanishing gradients, a novel angle-optimized text embedding model called AnglE is proposed by introducing angle difference optimization in a complex space which substantially improve the text embedding quality in various scenarios. Given input text embedding pair $(E(q),E(d))$, the chunking strategy [84] is used to get their representations in the complex space $(\mathbf{z},\mathbf{w})$, followed by the angle difference $\Delta\theta_{qd}$ between $\mathbf{z}$ and $\mathbf{w}$. Then the angle loss can be defined as:
@@
where $sim(E(i),E(j))$ is the similarity between the embedding of $i$ and the embedding of $j$. The authors also propose LLM-supervised learning (use LLMs as data annotators to label the pseudo-supervised data) to effectively deal with the domain-supervised data scarcity problem [50]. Some other details about UAE include:
  • Training data: the NLI datasets MNLI [73] and SNLI [74], and/or LLM supervised data
@@
Matryoshka Representation Learning (MRL) [62]
Deploying deep representation or text embedding involves two steps: a constant forward pass to compute the representation, and its use for downstream applications [85, 86]. The computation costs for the second step rise with the embedding dimensionality, data size, and label space, which can exceed the feature computation cost for large scale systems [87, 88]. The rigid nature of these representations requires high-dimensional embedding vectors for different tasks, even though varying resource and accuracy constraints call for flexibility [62]. Given that we can’t predict the computational and statistical demands for each subsequent task, fixed-capacity representations/embeddings may not always be suitable and could either exceed or fall short of the task’s requirements. Could we create an adaptable representation that can adjust to a variety of downstream tasks with fluctuating computational resources?
MRL introduces a novel method for learning representations of data through a nested structure to induce flexibility in the learned representation, similar to Russian Matryoshka dolls, which encodes information at different granularities and allows a single embedding to adapt to the computational constraints of downstream tasks [62]. The representation/embedding $z$ is a $d$ dimensional vector, $M=[m_{1},m_{2},...d]$ are the chosen dimensions which define different representation sizes. MRL makes each of the first $m$ dimensions $z_{1:m}$ to be independently capable of being a general purpose representation of the data point $x$. Given a labelled dataset $D=\{(x_{1},y_{1}),(x_{2},y_{2}),...(x_{N},y_{N})\}$ where $N$ is the datasize and $y_{i}$ is the label of data $x_{i}$, MRL uses standard empirical risk minimization to optimize multi-class classification loss for each nested dimension $m\in M$ using a separate linear classifier, parameterized by $\mathbf{W}^{(m)}$:
@@
where $\mathcal{L}$ is the multi-class softmax cross-entropy loss function, $F(\cdot;\theta_{F})$ is the deep neural network to get the representation/embedding $z$, $c_{m}$ is the importance scales. The authors also show that MRL extends seamlessly to web-scale datasets across vision, language, and vision + language. The experimental results show that MRL can be effectively used for large-scale adaptive classification and retrieval, providing similar accuracy to fixed-feature baseline with a significantly smaller representation size, and offering a more cost-effective and faster adaptive shortlisting and re-ranking system [62].
@@
Despite MRL’s enhanced efficiency, it still necessitates going through all transformer layers before obtaining the transformer based text embedding, leading to significant compute and memory consumption. This raises questions about the impact of the fixed number of transformer layers on representation quality and the feasibility of using intermediate layers for sentence representation. With the aim to enhance the flexibility and scalability of the original MRL’s sentence embedding learning, two-dimensional Matryoshka Sentence Embedding (2DMSE) is proposed in [63]. 2DMSE uses $BERT_{base}$ as backbone to encode text data $x$:
@@
The auxiliary information A is utilized for loss computation, typically indicating positive or negative samples or providing ranking details [63]. During the same training step, a shallower Transformer layer $l$ is randomly chosen following a uniform distribution $l\sim U(1,L-1)$, and its complete embedding vector is directly utilized for representation learning:
@@
Based on MRL and 2DMSE, several well performing text embeddings including mxbai-embed-large-v1 (335M) and mxbai-embed-2d-large-v1 (335M) are released in [89]. However, the training details of these models are not documented.